# MCCoder: Streamlining Motion Control with LLM-Assisted Code Generation and Rigorous Verification

### v1

@@ V-E Error Analysis
    API errors (5.2%): These errors indicate incorrect API function invoking. But the API error rate is relatively small, showing the effectiveness of the current retrieval method for relevant sample codes reference.
@@
    Argument errors (79%): The most prevalent error type involves correct API invocations with incorrect arguments, frequently stemming from hallucinations or disregard for prompt instructions. Self-correction can sometimes address and correct these errors. Explicit error messages like ”type object ProfileType has no attribute scurve. Did you mean: SCurve?” significantly enhance the likelihood of successful correction in subsequent attempts. In contrast, implicit error messages hinder accurate identification and correction.
@@
    Syntax errors (15.8%): These errors predominantly arise from hallucinations or the invention of nonexistent functions and libraries, resulting in syntactic inaccuracies.
V-F Ablation Study
We conducted an ablation study focusing on prompt engineering, task decomposition, and self-correction modules with the base models GPT-4o and GPT-4o-mini. As shown in Figure 4, prompt engineering plays a significant role in the MCCoder system, causing a 20.51% performance drop in FTPR (DTW) for GPT-4o and a 15.79% drop for GPT-4o-mini, with a larger impact on the larger model. Task decomposition resulted in a performance drop of around 13% for both models, indicating that more than half of the difficulty level 3 tasks failed without this module. The self-correction module had a relatively smaller impact, with a 1-5% performance difference. However, as we improve the precision of the soft-motion controller error information in the future, this module is expected to play a more substantial role.
Figure 4: Ablation study of MCCoder. Prompt engineering and task decomposition play critical roles in the system, while self-correction module shows a relatively small impact because of accuracy of the error information.
@@ VI DISCUSSION
VI-A Engineers’ feedback
We presented the system to a group of control engineers, ranging from novices to experts, and collected their feedback. Both groups highly regarded MCCoder for facilitating learning new functions and improving coding efficiency. They highlighted the necessity of human-involved verification to address safety concerns, an issue they encountered when using general LLMs for control code generation. Moreover, they suggested enhancing code interpretability, such as through visual flowcharts or explanations of API usage, to assist in better understanding and debugging when minor errors occur. This feedback has provided valuable insights for future improvements and deployment.
VI-B Limitations
The current MCEVAL dataset may not comprehensively encompass the full spectrum of motion control tasks, necessitating further testing across a broader range of motion scenarios to enhance the dataset. Additionally, the error messages returned by the soft-motion controller are occasionally ambiguous, which impedes MCCoder’s self-correction capabilities. While the system’s effectiveness has been demonstrated through the four existing model evaluations, conducting additional model tests could yield further insights for refinement and improvement.
@@ VIII CONCLUSION
In this paper, we introduce MCCoder to leverage LLMs for generating motion control code. MCCoder addresses the complexities and safety-critical aspects of motion control programming by integrating prompt engineering, task decomposition, retrieval, code generation, soft-motion and verification through simulation and data logging. Validated using the MCEVAL dataset, MCCoder significantly improves code generation performance, especially for the most challenging tasks. These advancement provide valuable insights and guidelines for future deployment and research in industrial automation and control code generation. Future work will focus on expanding the MCEVAL dataset, improving error message clarity from the soft-motion controller, and exploring fine-tuning with larger models to further enhance MCCoder’s capabilities.
@@ ACKNOWLEDGMENT
This work is funded by National Natural Science Foundation of China Grant No. 72371271, the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), the Nansha Key Area Science and Technology Project (No. 2023ZD003), and Project No. 2021JC02X191.

### v2

@@ V-E Error Analysis
    API errors (8.2%): These errors indicate incorrect API function invoking. But the API error rate is relatively small, showing the effectiveness of the current retrieval method for relevant sample codes reference.
@@
    Argument errors (76.3%): The most prevalent error type involves correct API invocations with incorrect arguments, frequently stemming from hallucinations or disregard for prompt instructions. Self-correction can sometimes address and correct these errors. Explicit error messages like “ProfileType has no attribute s curve. Did you mean: SCurve?” significantly enhance the likelihood of successful correction in subsequent attempts. In contrast, implicit error messages hinder accurate identification and correction.
@@
    Syntax errors (15.5%): These errors predominantly arise from hallucinations or the invention of nonexistent functions and libraries, resulting in syntactic inaccuracies.
@@ VI DISCUSSION
VI-A Limitations and Future Work
While MCCoder’s overall architecture and design can be applied to programming motion control tasks across various motion libraries using high-level languages, its current validation has been limited to the WMX3 library with Python. Future work will focus on expanding MCCoder’s applicability by collaborating with more motion control manufacturers and incorporating additional programming languages such as C++ and C#. This will enhance MCCoder’s versatility and extend the applicability of MCEVAL’s evaluation data across a broader range of motion control systems.
The broader adoption of AI-assisted systems in industrial applications requires not only advancements in AI itself but also improvements in controller components. A key limitation is the need for more detailed error reporting and real-time interfaces, which are crucial for enhancing system efficiency and robustness.
Furthermore, deploying MCCoder in real-world production environments necessitates fine-tuned, edge-side small models tailored to proprietary motion libraries. Integrating these specialized models with RAG techniques will significantly enhance data security and mitigate model hallucinations, making this an important research direction moving forward.
@@ VIII CONCLUSION
In this paper, we introduce MCCoder to leverage LLMs for generating motion control code. MCCoder addresses the complexities and safety-critical aspects of motion control programming by integrating prompt engineering, task decomposition, retrieval, code generation, soft-motion and verification through simulation and data logging. Validated using the MCEVAL dataset, MCCoder significantly improves code generation performance, especially for the most challenging tasks. These advancement provide valuable insights and guidelines for future deployment and research in industrial automation and control code generation. Future work will focus on enhancing MCCoder’s versatility, improving error message clarity from the soft-motion controller, and exploring fine-tuning with edge-side small models to further enhance its capabilities and practicability.
@@ ACKNOWLEDGMENT
This work is funded by National Natural Science Foundation of China Grant No. 72371217, the Guangzhou Industrial Informatic and Intelligence Key Laboratory No. 2024A03J0628, the Nansha Key Area Science and Technology Project No. 2023ZD003, and Project No. 2021JC02X191.